OpenStudy (anonymous):

Maria solved the equation -5 1/8 n = -9 3/4. Her answer was n = 1 37/41. She is fairly confident about her answer, but wants to do a quick check of her solution using estimation. Which of the following shows a good check of her answer?

OpenStudy (math&ing001):

You just need to plug the value of n in your equation like this : -5 1/8 * 1 37/41 = -9 3/4 and see if it checks

OpenStudy (anonymous):

these are my options A. -6(2) = -12 B. -5(-2) = -10 C. -5(2) = -9 D. -5(2) = -10

OpenStudy (math&ing001):

Estimation means you round up to the nearest hole number. In our case : -5 1/8 ~ -5 and 1 37/41 ~ 2 and -9 3/4 ~ -10 So what option should it be ?

OpenStudy (anonymous):

i dont know..... thats why i asked...

OpenStudy (math&ing001):

I gave you all hints so that all you have to do is replace. But anyways.. From (-5 1/8) * (1 37/41) = (-9 3/4) And -5 1/8 ~ -5 and 1 37/41 ~ 2 and -9 3/4 ~ -10 We get -5 * 2 = -10